# Explainable AI for Autism Diagnosis: Identifying Critical Brain Regions Using fMRI Data

### v1

@@ Abstract
Early diagnosis and intervention for Autism Spectrum Disorder (ASD) has been shown to significantly improve the quality of life of autistic individuals. However, diagnostics methods for ASD rely on assessments based on clinical presentation that are prone to bias and can be challenging to arrive at an early diagnosis. There is a need for objective biomarkers of ASD which can help improve diagnostic accuracy. Deep learning (DL) has achieved outstanding performance in diagnosing diseases and conditions from medical imaging data. Extensive research has been conducted on creating models that classify ASD using resting-state functional Magnetic Resonance Imaging (fMRI) data. However, existing models lack interpretability. This research aims to improve the accuracy and interpretability of ASD diagnosis by creating a DL model that can not only accurately classify ASD but also provide explainable insights into its working. The dataset used is a preprocessed version of the Autism Brain Imaging Data Exchange (ABIDE) with 884 samples. Our findings show a model that can accurately classify ASD and highlight critical brain regions differing between ASD and typical controls, with potential implications for early diagnosis and understanding of the neural basis of ASD. These findings are validated by studies in the literature that use different datasets and modalities, confirming that the model actually learned characteristics of ASD and not just the dataset. This study advances the field of explainable AI in medical imaging by providing a robust and interpretable model, thereby contributing to a future with objective and reliable ASD diagnostics.
@@
Biomarkers, Deep Learning, Interpretability, Sparse Autoencoders, RemOve And Retrain (ROAR)
@@ I Introduction
Autism Spectrum Disorder (ASD) is a complex neurodevelopmental condition characterised by impairments in social communication, interaction, and restricted, repetitive behaviors [1]. Its presentation varies widely across affected individuals, impacting on quality of life. In the last 20 years, there has been a 787% surge in ASD diagnosis increasing its prevalence from approximately 1% to around 3% of the population. This surge in diagnosis is thought to be due to the delay in recognition of the clinical symptoms and the lack of personnel to diagnose. [2, 3, 4]. Early diagnosis, ideally between the ages of two and five, can lead to a significantly better Quality of Life (QoL) for individuals with ASD [5]. Studies also show that early intervention can improve IQ scores by up to 18 points and enhance language skills [6], emphasising the critical need for early diagnosis and intervention. However, the average age of diagnosis remains around five years old, highlighting a crucial limitation in current diagnostic practices [7].
@@
Specifically, current clinical ASD diagnostic practices lack effective early diagnosis methods. The method of diagnosing ASD in children is behavioral analysis combined with clinical history, which is not effective for early diagnosis as it is based on identifying abnormal communication and social behaviors that may not emerge until the condition is well established [8]. It is also subject to clinicians’ biases and has historically led to under-diagnosing ASD particularly, in certain groups such as females [9]. To enhance our understanding of ASD and improve diagnostic accuracy, there is a need for objective, biologically-based markers of the condition.
@@
The resting-state functional Magnetic Resonance Imaging (fMRI) data of people with ASD has been analysed to show abnormalities in functional connectivity [^1] Functional connectivity examines the correlation of activity between different brain regions. If two areas consistently show increases and decreases in activity together, they are considered functionally connected. [10]. These differences in distinctive patterns in the functional connectivity of autistic individuals can serve as biomarkers of ASD. By researching these differences, studies have identified several brain networks and areas that are associated with ASD such as disrupted connectivity in the Default Mode Network (DMN) [11]. However, the vast differences in ASD and the concept of the “idiosyncratic brain” suggests that there exists a highly personalised neural network configuration for each individual [12, 10]. The regions involved are also affected by the severity of ASD [13]. This highlights the need for a method to identify the intricate patterns of neural networks in ASD and possibly reveal a biomarker that is characteristic of all ASD individuals regardless of the severity of ASD.
@@
These abnormalities can be learned using deep learning (DL) to distinguish between ASD and Typical Control (TC). Many recent studies have emerged using functional connectivity for ASD classification using machine learning, with some showing a high accuracy ($>$ 90%) [14, 15]. While these models cannot replace clinical professionals, they can offer valuable decision support. These neuroimaging-based diagnosis methods have the potential to revolutionise ASD diagnosis, but most current approaches lack in explainability. These limitations reduce trust and hinder clinical acceptance of the model in real-world scenarios [16]. To address this challenge, there is a large body of research focusing on approaches that enhance the interpretability of DL models in the medical domain [17]. However, identifying which approach should be used with functional connectivity data or how they should specifically be used to identify biomarkers in ASD has not yet been explored.
@@
This study explores the utilisation of DL methods to identify potential biomarkers for ASD through the analysis of resting-state fMRI data, aiming to develop and explain a predictive model that is both accurate and interpretable. We also benchmark a variety of interpretability methods using the Remove And Retrain (ROAR) technique [18] to identify the most effective method for finding important regions in the brain for ASD classification.
@@
Successful application of the outcomes of this research could lead to improved access to early interventions, reduced costs of ASD diagnosis, and the development of more effective diagnostic tools. By enhancing our understanding of the neurobiological markers of ASD, this work aims to contribute to the development of more targeted interventions and improved quality of life for individuals with ASD and their families.
@@ II Related Work
II-A Neurological basis of ASD and fMRI studies
Research into ASD using fMRI has revealed distinct connectivity patterns between individuals with ASD and typical controls. Understanding these patterns is vital as they provide objective, neurobiological markers that could enhance early detection and inform targeted interventions. Functional connectivity, which measures synchronised activity between different brain regions, shows atypicality in several areas frequently linked to ASD. Key regions include the frontal cortex, involved in executive functions and social cognition [20], temporal regions for language and social processing [21], the amygdala for emotion processing [22], and the cerebellum for motor function and social cognition [23]. Studies indicate that ASD individuals often exhibit local over-connectivity and long-range under-connectivity [24, 25, 26]. Furthermore, disrupted connectivity in the DMN and Salience Network (SN) is associated with ASD [11, 27]. However, there is considerable variation in the important regions identified across studies, and these regions can change on a case-by-case basis depending on the severity of ASD symptoms [13]. Notably, a universal biomarker common to all cases of Autism, if it exists, has not been found. The concept of the “idiosyncratic brain” suggests highly personalised neural network configurations in ASD, presenting a complex challenge for traditional analysis methods. This emphasises the need for advanced techniques like DL [^2] Traditional ML methods struggle with the high dimensionality and complex, non-linear relationships present in neuroimaging data. Deep learning, with its ability to automatically learn hierarchical representations from raw data, is particularly well-suited for capturing the subtle and highly variable connectivity patterns observed in ASD. to identify the intricate patterns in ASD neuroimaging data [12].
II-B Deep Learning for brain imaging
DL techniques, inspired by the structure of the brain, have proven highly effective in medical imaging, surpassing traditional machine learning methods [28]. DL models capture complex patterns and combine data from multiple modalities, improving diagnostic accuracy. In brain imaging, various DL architectures like Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), and Autoencoders are employed for different analysis tasks [29, 30, 31]. Challenges include managing high-dimensional and multimodal data, mitigating overfitting, and ensuring model interpretability. Initiatives promoting data sharing, ethical frameworks, and cloud-based resources are addressing these challenges, while explainable AI research is gaining traction to identify key brain regions driving DL model predictions. Ethical considerations are paramount, as DL models must be evaluated rigorously to prevent biases and discriminatory practices [32]. In the context of ASD diagnosis, these advanced DL techniques show great promise, offering the potential for high accuracy and early detection by leveraging complex patterns in brain imaging data that may not be apparent to the human eye or traditional analysis methods.
@@ II-C Feature analysis and interpretability in ASD research
Feature analysis helps translate a model’s accuracy into insights about the underlying mechanisms it uses for classification. Methods include gradient-based techniques like Layer-wise Relevance Propagation (LRP) [33] and DeepLift [34], and perturbation techniques like SHapley Additive exPlanations (SHAP)[35] and Local Interpretable Model-Agnostic Explanations (LIME) [36]. In medical research, these interpretability methods have been successfully applied to identify biomarkers for various conditions [37, 38, 39]. However, a comprehensive review [40] found that there is no universally “best” interpretability method; rather, their effectiveness varies depending on the specific application and data type.
In the context of ASD research, there is a notable gap in the application of these interpretability methods, particularly for functional connectivity data derived from fMRI. This gap is exacerbated by several challenges unique to ASD and neuroimaging data. Firstly, the heterogeneity of ASD manifestations adds another layer of complexity, as different neural patterns may be significant for different subgroups within the ASD spectrum. Secondly, the lack of ground truth for important brain regions in ASD makes it challenging to validate the results of interpretability methods. Moreover, even when studies attempt to incorporate interpretability, they often fail to validate their results, leaving uncertainty about whether the model is truly learning characteristics of Autism or merely recognising patterns specific to the dataset. While some studies have attempted to use similar approaches in related fields, such as the use of back-propagation to identify biomarkers for Attention Deficit Hyperactivity Disorder (ADHD) [41], these efforts often suffer from limitations like low accuracy and lack of proper benchmarking. The absence of a standardised framework for evaluating and comparing different interpretability methods in the context of ASD further complicates the situation. These challenges collectively highlight the critical need for a systematic approach to evaluate and apply interpretability methods in ASD research, which is precisely what this study aims to address.
@@ II-D Gaps in current research
A significant challenge in ASD research is the lack of reliable biomarkers, which hinders early diagnosis. Despite numerous studies, a universal biomarker common to all cases of Autism has not been identified, largely due to the heterogeneous nature of ASD and the notion of unique neurological variations among individuals [12]. This absence of definitive neurobiological markers significantly impedes the development of objective, early diagnostic tools for ASD.
Many studies in ASD diagnosis using DL face methodological challenges. Notably, there’s often a limited focus on evaluation metrics, with many studies reporting only accuracy, ignoring other crucial metrics like sensitivity and specificity [42]. This narrow focus can lead to overfitting and a false sense of model performance. Moreover, while some studies have achieved high accuracy, there’s a lack of explainable models with accuracy exceeding 90%. This gap raises a critical question: how can we trust an explainable model if it’s not sufficiently accurate, and conversely, how can we rely on a highly accurate model if we can’t explain its decisions?
The majority of studies have not employed interpretability methods to elucidate the regions utilised by their models. This results in “black box” models that may achieve high performance but lack a mechanism to verify whether they are genuinely learning to identify Autism or merely memorising dataset-specific patterns. Furthermore, when interpretability methods are used, they are often not benchmarked against each other, leaving uncertainty about which methods are most reliable for ASD research.
Another critical issue is the lack of diverse datasets representing various ethnicities, socioeconomic backgrounds, and ASD severity levels, which limits the models’ ability to generalise across diverse populations [43]. This could lead to biased diagnoses and exacerbate existing health inequalities if implemented in clinical settings without proper validation.
This project aims to address these critical gaps by developing an explainable DL model for ASD classification with high accuracy, rigorously benchmarking interpretability methods, and identifying key brain regions contributing to the diagnosis. While addressing the diversity issue is crucial for future clinical applications, it is beyond the scope of this research due to limited resources and the constraints of available datasets. Nevertheless, our study paves the way for more reliable, transparent, and potentially generalisable AI-assisted ASD diagnostics.
@@ III-A1 Dataset
This study used the Autism Brain Imaging Data Exchange (ABIDE) dataset [44], specifically the Preprocessed Connectomes Project (PCP) version. ABIDE was chosen for its large sample size, ease of access, and prevalence in comparable studies, facilitating performance benchmarking.
The dataset includes resting-state fMRI data from 539 individuals with ASD and 573 typically developing controls, aged 7-64 years. While ABIDE offers a robust sample size for model training, it has a notable gender bias (85.2% male) [45]. Notably, We employed stratified k-fold cross-validation during model training and evaluation, ensuring that each fold maintained the original dataset’s distribution. We acknowledge that our results should be interpreted with caution regarding generalisability across genders. The data is anonymised in accordance with HIPAA guidelines and 1000 FCP / INDI protocols, ensuring ethical use of the neuroimaging data.
@@ III-A2 Preprocessing
The PCP openly shares preprocessed neuroimaging data from ABIDE. There is no set standard on how neuroimaging data should be preprocessed so PCP offers data preprocessed using four different preprocessing pipelines, namely the Connectome Computation System (CCS) [46], the Configurable Pipeline for the Analysis of Connectomes (CPAC) [47], the Data Processing Assistant for Resting-State fMRI (DPARSF) [48] and the NeuroImaging Analysis Kit (NIAK) [49]. The preprocessing steps carried out by the various pipelines are comparatively analogous [14]. The main differences are in the particular algorithms for each step, the software simulations, and the parameters applied. Details of each pipeline technique are provided on the PCP website [50]. Since there is no consensus on what precprocessing pipeline is the best we have trained a model using each of the four pipelines and compared their results.
PCP applies some standard low-level fMRI preprocessing steps to all preprocessing pipelines. These include brain extraction, spatial smoothing, temporal filtering, realignment (motion correction), slice timing correction and intensity normalisation [47]. Further, there has been significant discussion in the literature about whether Framewise Displacement (FD)  [^3] Framewise Displacement (FD) is a measure of how much head movement occurs during the MRI scan filtering affects the accuracy of an autism diagnosis model. This necessitates an investigation to clarify head movement’s role in model performance. Some influential studies state that an FD value $>$ 0.2mm corrupts the fMRI data [51, 52], leading to a decrease in performance. In contrast, another study stated that FD filtering or head movement did not affect the performance of their model [53]. By conducting our study under both conditions - with and without FD filtering - we aim to provide clarity on this methodological issue and its impact on model accuracy.
@@ III-A3 Parcellation
Brain images are overlaid on a brain atlas, which divides the brain into several Regions of Interest (ROIs). The voxels within each ROI are averaged to produce a single value, effectively reducing the dimensionality of the data. For this project, we selected the Automated Anatomical Labelling (AAL) atlas, which offers a balanced approach with 116 ROIs. This choice was driven by computational efficiency and proven performance. The computational cost of our model increases quadratically with the number of ROIs due to the use of functional connectivity matrices (FCM, explained in III.B), and the AAL atlas provides a good balance between granularity and manageable computational demands. Moreover, the AAL atlas has demonstrated high accuracy in similar studies [54, 15], supporting its suitability for ASD classification tasks. After preprocessing and atlas application, our dataset comprises of 884 samples (408 ASD, 476 TC). Each sample is represented as a 2-dimensional array $(N\texttimes t)$, where $t$ is the temporal dimension and $N$ corresponds to the number of ROIs in our atlas (116 in our case).

### v2

@@ Abstract
Early diagnosis and intervention for Autism Spectrum Disorder (ASD) has been shown to significantly improve the quality of life of autistic individuals. However, current diagnostic practices depend on behavioral assessments, which are prone to bias and fail to capture early manifestations of the disorder, making early diagnosis challenging. There is a need for objective biomarkers of ASD which can help improve early diagnostic accuracy. Extensive research has been conducted on creating machine learning models that classify ASD using resting-state functional Magnetic Resonance Imaging (fMRI) data. However, existing models lack interpretability. This research aims to improve the accuracy and interpretability of ASD diagnosis by creating a deep learning (DL) model that can not only accurately classify ASD but also provide explainable insights into its working. The dataset used is a preprocessed version of the Autism Brain Imaging Data Exchange (ABIDE). Our findings show a model that achieved a classification accuracy of 98.2% and an F1 score of 0.97, while also highlighting critical brain regions differing between ASD and typical controls, with implications for early diagnosis and understanding of the neural basis of ASD. Additionally, a systematic evaluation of interpretability methods using the Remove And Retrain (ROAR) technique highlights the reliability of gradient-based approaches for fMRI data. These findings are validated by studies in the literature that use different datasets and modalities, confirming that the model actually learned characteristics of ASD and not just dataset artifacts.
@@
Biomarkers, Deep Learning, Explainable AI, Functional Connectivity, Functional MRI (fMRI), Interpretability, Magnetic Resonance Imaging, Neuroimaging, RemOve And Retrain (ROAR), Sparse Autoencoders
@@ I Introduction
Autism Spectrum Disorder (ASD) is a complex neurodevelopmental condition characterised by impairments in social communication, interaction, and restricted, repetitive behaviors [1]. Its presentation varies widely across affected individuals, impacting on quality of life. In the last 20 years, there has been a 787% surge in ASD diagnosis increasing its prevalence from approximately 1% to around 3% of the population. This surge in diagnosis is thought to be due to the delay in recognition of the clinical symptoms and the lack of personnel to diagnose [2, 3, 4]. Early diagnosis, ideally between the ages of two and five, can lead to a significantly better Quality of Life (QoL) for individuals with ASD [5]. Studies also show that early intervention can improve IQ scores by up to 18 points and enhance language skills [6], emphasising the critical need for early diagnosis and intervention. However, the average age of diagnosis remains around five years old, highlighting a crucial limitation in current diagnostic practices [7].
@@
Specifically, current clinical ASD diagnostic practices lack effective early diagnosis methods. The method of diagnosing ASD in children uses behavioral analysis combined with clinical history, which is not effective for early diagnosis as it is based on identifying abnormal communication and social behaviors that may not emerge until the condition is well established [8]. It is also subject to clinicians’ biases and has historically led to under-diagnosing ASD particularly, in certain groups such as females [9]. To enhance our understanding of ASD and improve diagnostic accuracy, there is a need for objective, biologically-based markers of the condition.
@@
The resting-state functional Magnetic Resonance Imaging (rs-fMRI) data of people with ASD has been analysed to show abnormalities in functional connectivity [^1] Functional connectivity examines the correlation of activity between different brain regions. If two areas consistently show increases and decreases in activity together, they are considered functionally connected. [10]. These differences in distinctive patterns in the functional connectivity of autistic individuals can serve as biomarkers of ASD. However, recent research highlights the “idiosyncratic brain” concept, where functional connectivity patterns vary widely among individuals with ASD. This variability complicates the search for a universal biomarker and underscores the need for a tool that can analyse these subtle differences. (For additional detail on these neural connectivity patterns, see Section II-A.)
@@
These differences can be learned using deep learning (DL) and be used to distinguish ASD and Typical Control (TC). Many recent studies have emerged using functional connectivity for ASD classification using machine learning, with some showing a high accuracy ($>$ 90%) [11, 12]. While these models cannot replace clinical professionals, they can offer valuable decision support, but most current approaches lack in explainability. These limitations reduce trust and hinder clinical acceptance of the model in real-world scenarios [13]. To address this challenge, there is a large body of research focusing on approaches that enhance the interpretability of DL models in the medical domain [14]. However, identifying which approach should be used with functional connectivity data or how they should specifically be used to identify biomarkers in ASD has not yet been explored.
@@
This study explores the utilisation of DL methods to identify potential biomarkers for ASD through the analysis of resting-state fMRI data, aiming to develop and explain a predictive model that is both accurate and interpretable. We also benchmark a variety of interpretability methods using the Remove And Retrain (ROAR) technique [15] to identify the most effective method for interpreting DL models that are trained on functional connectivity fMRI data.
@@
Successful application of the outcomes of this research could lead to improved access to early interventions, reduced costs of ASD diagnosis, and the development of more effective diagnostic tools. By enhancing our understanding of the neurobiological markers of ASD, our work aims to contribute to the development of more targeted interventions and improved quality of life for individuals with ASD and their families.
@@ II Related Work
II-A Neurological Basis of ASD
Autism Spectrum Disorder (ASD) is associated with distinct yet heterogeneous patterns of functional connectivity in the brain, often manifesting as local over-connectivity and long-range under-connectivity [17, 18, 19]. Key regions implicated include the frontal cortex (executive functions, social cognition) [20], temporal cortex (language processing) [21], amygdala (emotion processing) [22], and the cerebellum (motor control, social cognition) [23]. Abnormalities in large-scale networks, such as the Default Mode Network (DMN) and Salience Network (SN), are also linked to ASD [24, 25]. However, considerable variability arises from the “idiosyncratic brain” concept [26], wherein each individual with ASD may exhibit unique functional connectivity patterns, making the search for a universal biomarker challenging. Moreover, symptom severity further modulates which regions display the greatest atypicality [27]. These factors collectively underscore the complexity of ASD’s neurobiological underpinnings and highlight the need for sophisticated methods—particularly deep learning (DL)—that can detect subtle and individualised patterns.
II-B Deep Learning Approaches for ASD
Deep learning has emerged as a promising tool in ASD research due to its ability to extract complex, nonlinear features from high-dimensional data such as fMRI [28, 29, 30]. Various architectures (e.g., CNNs, RNNs, GNNs, and autoencoders) have been explored to improve diagnostic accuracy and potentially enable early detection [29, 30, 31]. However, existing studies vary widely in sample size, accuracy, and interpretation strategies.
For example, a study [32] applied a deep neural network classifier to a small dataset of fewer than 130 samples, achieving accuracies between 85.3% and 87.1%. Although they introduced a single interpretability method, they did not determine whether the highlighted features were consistent with known ASD-related neurobiology. In a separate study, [12] leveraged over 1,000 samples from the Autism Brain Imaging Data Exchange (ABIDE) to attain a high accuracy of 93.59% using a Stacked Sparse Autoencoder (SSAE), but offered no mechanism to clarify which specific brain connections contributed most to classification. Independently, another group of authors [33] focused on biomarker discovery and employed an interpretability method, yet their accuracy was lower at around 72% and they did not cross-reference the discovered biomarkers with established ASD findings. Collectively, these studies demonstrate the potential of DL in ASD classification, particularly for large-scale data, but also reveal critical shortcomings: either interpretability is absent, or when present, accuracy is low and it is rarely corroborated with well-established ASD findings from the neuroscience literature.
@@ II-C Feature Analysis and Interpretability in ASD Research
Interpretability methods bridge the gap between raw classification accuracy and meaningful clinical insights. Gradient-based algorithms, such as Layer-wise Relevance Propagation (LRP) [34] and DeepLift [35], and perturbation-based tools like SHAP [36] or LIME [37], have been used successfully in other medical contexts [38, 39, 40]. Yet, ASD research presents unique challenges: the heterogeneous presentation of ASD may mean that each subgroup has distinct connectivity markers, there is no absolute ground truth for “crucial” regions in ASD, and most existing studies that do employ interpretability methods fail to confirm whether their highlighted regions align with validated ASD neurobiology [41, 42]. Further compounding this is the lack of a standard framework for benchmarking interpretability in ASD-specific fMRI data, leaving researchers to rely on a single method without broader validation or comparisons.
@@ II-D Gaps in Current Research
While deep learning models can offer high accuracy in ASD classification, several persistent gaps limit both clinical utility and scientific understanding:
  1. 1.
    Lack of Transparent DL Models: Many high-accuracy studies [12] do not explain which brain regions or connections drive their decisions, leading to “black box” models.
  2. 2.
    Limited Interpretability Benchmarking: The few studies that do focus on interpretability only use a single approach to highlight features, leaving uncertainty about whether alternate methods would reveal or confirm the same features.
  3. 3.
    Sparse Validation Against Neuroscience: Even when interpretability is attempted, few works cross-reference identified biomarkers with established ASD findings, raising doubts about clinical relevance [32, 33].
These challenges highlight the need for an approach that combines high diagnostic performance with robust interpretability and neuroscientific validation. In this paper, we address these gaps by:
  • Developing a DL pipeline that can classify and interpret ASD with high accuracy.
  • Evaluating multiple interpretability methods in a systematic benchmark.
  • Cross-referencing the identified brain regions with established ASD research to ensure clinical and biological relevance.
By uniting rigorous performance evaluation, interpretability, and domain validation, we address the critical need for a transparent system that can unveil meaningful biomarkers and build clinical trust.
@@ III-A1 Dataset
This study used the Autism Brain Imaging Data Exchange (ABIDE) dataset [43], which comprises resting-state fMRI data from 539 autistic and 573 typically developing individuals aged 7–64 years from 16 international imaging sites. Note that the sample’s gender imbalance (85.2% male) may limit generalisability [44]. The data is anonymised in accordance with standard ethical protocols and were made available under the 1000 Functional Connectomes Project (FCP) and International Neuroimaging Data-Sharing Initiative (INDI).
@@ III-A2 Preprocessing
The Preprocessed Connectomes Project (PCP) openly shares ABIDE data preprocessed via four pipelines—the Connectome Computation System (CCS) [45], the Configurable Pipeline for the Analysis of Connectomes (CPAC) [46], the Data Processing Assistant for Resting-State fMRI (DPARSF) [47], and the NeuroImaging Analysis Kit (NIAK) [48]. These pipelines are intended to mitigate center-specific artifacts. Because no single pipeline is universally optimal, we trained our model on all four to compare performance [49]. Additionally, analysing data from multiple pipelines helps us identify the core regions that remain consistent regardless of preprocessing choices, ensuring that any discovered biomarkers are pipeline-independent. Full details on these pipelines are available in the PCP documentation [49].
Each pipeline applies standard low-level preprocessing steps, such as brain extraction, realignment (motion correction), and slice timing correction [46]. Further, there has been significant debate in the literature about whether Framewise Displacement (FD) [^2] Framewise Displacement (FD) is a measure of how much head movement occurs during the MRI scan filtering influences ASD classification performance. Some influential studies state that an FD value >0.2 mm corrupts the fMRI data [50, 51], leading to decreased accuracy, whereas another study [52] contends that motion does not appreciably affect performance. We conducted experiments both with and without FD filtering to clarify this methodological issue.
Applying an FD threshold of 0.2 mm yielded 884 usable samples (408 ASD, 476 TC). These were drawn from multiple international sites and scanners, which increases data diversity but may introduce additional variability.
@@ III-A3 Parcellation
Brain images were then parcellated with the Automated Anatomical Labelling (AAL) atlas, which divides the brain into 116 Regions of Interest (ROIs). We selected the AAL atlas for its balance of anatomical detail and manageable complexity, as recommended in similar ASD studies [12, 53]. This produced an $N\times t$ data array (where $N=116$ ROIs, $t$ is the temporal dimension) for each of the 884 subjects, providing the foundation for subsequent feature extraction.